OpenStudy (anonymous):

i couldn't find the answer for this question, not sure if i'm remotely close the problem says solve cos 2x=2(sinx-1) for all real x.

OpenStudy (anonymous):

cos 2x can be expressed in terms of sin^2(x), right? If you do that, you get a quadratic in sin (x) which can be solved. Try it!

OpenStudy (anonymous):

so next step is sin^2 (x) = 2sinx - 2 ?

OpenStudy (anonymous):

How do we express cos(a+b)?

OpenStudy (anonymous):

cos(a+b)=cos(a).cos(b)-sin(a).sin(b), remember?

OpenStudy (anonymous):

wait i remember that, but cos2x in the beginning can turn into 1-2sin^2? right

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

So if you think of sin^2(x) as some variable, say y, we get a quadratic equation in y which can be solved.

OpenStudy (anonymous):

1-2sin^2(x)= 2(sinx - 1)

OpenStudy (anonymous):

Yes right.

OpenStudy (anonymous):

so then i try to express it in a way that it looks like a^2 + a value - a value and then factor?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

not sure if this is right, 2sin^2 (x) -2sinx-1=0

OpenStudy (anonymous):

Let me check

OpenStudy (anonymous):

Lets work it out: \[\cos(2x)=1-2\sin^{2}(x)\] So: \[1-2\sin ^{2}(x)=2(\sin(x)-1)=2\sin(x)-2\]

OpenStudy (anonymous):

Gives me: \[2\sin ^{2}(x)+2\sin(x)-3=0\]

OpenStudy (anonymous):

Right??

OpenStudy (anonymous):

oh, i see where i went wrong

OpenStudy (anonymous):

Now on it is the application of the quadratic formula to get the value of sin(x) and hence (x)

OpenStudy (anonymous):

okay, and then set them to zero meaning (first equation)=0 (second equation) = 0

OpenStudy (anonymous):

Not sure what you mean. I would substitute sin(x) by another variable, say y to get a quadratic as:\[2y ^{2}+2y-3=0\]

OpenStudy (anonymous):

The solve for y

OpenStudy (anonymous):

Got it?